# On the Age of Information in Single-Server Queues with Aged Updates

### v1

@@ 4.1.2 Light traffic
The LT regime occurs when the effective arrival rate approaches zero. Together with the Global Balance Principle, this feature is captured by the condition
@@
For a given standard deviation, this implies that the coefficient of variation $\kappa_{Y}$ goes to zero. Consequently, the second term in the right-hand side of equation (7) and the interval width in Corollary 1 become negligible, which leads to the independent case with very tight bounds.
@@ 4.2 Forwarding
To illustrate this, a zero-wait model is considered, in which the source is aware of the system’s state and can generate a fresh packet as soon as a delivery occurs. This model is idealised, but also insightful in capturing the best-case AoI performance [11, Section III.B]. In particular, because of the source responsiveness, the time variables in the zero-wait model are identically distributed,
$X_{n+1}=T_{n}=Y_{n}=S_{n}\sim\textsl{exp}\left(\mu\right).$ (9)
@@
Figure 3: Error-prone zero-wait model as a system with aged updates. a) Error-prone zero-wait standard model. Blue triangles are successful deliveries and generation of a fresh packet, and red triangles ($t_{3}$ and $t_{5}$) mark failed deliveries and subsequent retransmission of the same packet. b) Equivalent error-free zero-wait model with aged updates. Failed transmissions in the standard model can be thought of as inducing an independent process of initial ages $A_{n}$.
The value of the AAoI in a zero-wait model with error probability $\alpha$, as the one depicted in Figure 3a), is given by [11, Equation (22)],
$\Delta_{zw}=\dfrac{2}{\mu(1-\alpha)},$ (10)
where $\mu(1-\alpha)$ can be seen as the effective service rate. Notably, the same expression is obtained by applying Theorem 1 to the error-free model of Figure 3b) with an independent process of initial ages defined by
@@
where $\tilde{S}_{1}\sim\textsl{exp}\left(\mu(1-\alpha)\right)$ and $\tilde{S}_{2}\sim\Gamma\left(2,\mu(1-\alpha)\right)$ are implicitly derived from the number of retransmissions. From this fact and (9), it follows that $A_{n}$ is independent of any other variable. Thus, since $\mathbb{E}\left[A_{n}\right]=\frac{2\alpha}{\mu(1-\alpha)}$, Theorem 1 yields
$\Delta_{zw}^{A}$ $=\Delta_{zw}^{0}+\mathbb{E}\left[A_{n}\right]$
$=\dfrac{2}{\mu}+\dfrac{2\alpha}{\mu(1-\alpha)}=\dfrac{2}{\mu(1-\alpha)}.$
Figure 4: Contour plot for the width of the interval defined in Corollary 1 in the error-prone zero-wait model.
@@
$d=2\sigma_{A}=\dfrac{2\sqrt{2\alpha(2-\alpha)}}{\mu(1-\alpha)},$
resulting in more accurate approximations for models with lower error probabilities or higher service rates. Figure 4 confirms that, for a given value of $\mu$, the largest widths are reached for high values of $\alpha$.
@@ 4.3 Tandem Queues
4.3.1 Application to M/M/1$\infty$ $\rightarrow$ M/M/1$\infty$ tandem
First, the theorem that gives the AAoI at the end of this tandem is stated. While this result is already known, this example is useful to confirm that the approach with aged updates can also be accommodated in this model.
Theorem 2 ([14], Theorem 3111[14, Equation (13)] contains a typo.).
The AAoI at the end of two infinite-capacity queues in tandem (M/M/1$\infty$ $\rightarrow$ M/M/1$\infty$) with $\lambda$-Poisson arrivals, and i.i.d. exponential service times of rates $\gamma$ and $\mu$, is given by
@@ 4.3.1 Application to M/M/1$\infty$ $\rightarrow$ M/M/1$\infty$ tandem
Figure 5 shows a sample path of the age at the end of this tandem. Arrivals at the first queue, represented by black triangles, occur at epochs $t_{n}$. After waiting in the queue if the server is busy upon arrival, the updates are served in a $\gamma$-exponential time, depart from the first queue and join the second at epochs $t^{\prime}_{n}$, marked with inverted black triangles. If the second server is found busy, the update waits again until it starts service, where it spends a $\mu$-exponential time, leaving the system at epochs $t^{\prime\prime}_{n}$. The final departures are represented in Figure 5 by inverted blue triangles. Note that the designation of this tandem as M/M/1$\infty$ $\rightarrow$ M/M/1$\infty$ is completely justified, because it is well-known that the departure process from the first queue –and thus, the arrival process at the second queue– is also a $\lambda$-Poisson process.
Figure 5: Sample path for age in a tandem of two queues M/M/1$\infty$ $\rightarrow$ M/M/1$\infty$. Packets arrive at the system at epochs $t_{n}$, depart from the first queue and join the second at epochs $t^{\prime}_{n}$, and leave the system at epochs $t^{\prime\prime}_{n}$. Packets arriving at the second queue have a positive age equal to the system time in the first queue. Note that $Y_{n}^{(1)}$ are inter-departure times in the first queue and inter-arrival times in the second queue.
@@
where $Y_{n}^{(2)}$ is the inter-departure time in the second queue, $A_{n}$ is the initial age and $\Delta^{0}_{M/M/1/\infty}$ is the average age of M/M/1$\infty$-FCFS queue with zero-age updates. It was found in [2] (see Equation (17)), that
@@
Then, Theorem 2 follows by applying Equation 12, Lemma 1 and $\mathbb{E}\left[Y_{n}^{(2)}\right]=1/\lambda$ to (11).
@@
This procedure can also derive the AAoI expressions for other well-known tandems. Specifically, the independence between final inter-departure times and initial ages can be exploited in tandems consisting of two single-capacity queues, with both preemption in service [15, Equation (48)] and without preemption in service [14, Theorem 1].
4.3.2 $C+1$ M/M/1$\infty$ queues in tandem
As acknowledged in [14], while the rationale employed for the two-queued tandem of Section 4.3.1 can be extended for more than two queues, the practical computation of the AAoI in larger tandems involves a chain of conditional variables whose treatment is quite complex. This is where the result of Corollary 1 becomes particularly valuable, as it provides a straightforward interval for the AAoI of a tandem formed by any arbitrary number of M/M/1$\infty$ queues.
Suppose that the service rate of the $i$-th server is $\gamma_{i},\ i=1,\ldots,C+1$. For consistency with the previous section, the notation $\gamma_{C+1}=\mu$ is used interchangeably to denote the service rate of the last server. Conveniently, the notation $\Delta$ is used to denote the AAoI at the end of the tandem, and $\Delta_{i}$ to denote the AAoI of a single queue with service rate $\gamma_{i}$. The departure process at the end of such a tandem remains Poisson of rate $\lambda$, implying that the coefficient of variation of the inter-departure times is $\kappa_{Y}=1$. Additionally, the initial age of the updates entering the last queue is the sum of the system times in each of the preceding queues,
@@ 4.3.2 $C+1$ M/M/1$\infty$ queues in tandem
$\mathbb{E}\left[A_{n}\right]=\sum\limits_{i=1}^{C}\frac{1}{\gamma_{i}-\lambda}\quad\text{ and }\quad\sigma_{A}=\sqrt{\sum\limits_{i=1}^{C}\frac{1}{(\gamma_{i}-\lambda)^{2}}},\$
@@
Moreover, evidence suggests the AAoI at the end of this tandem is independent of the ordering of the servers [14]. Although this hypothesis has not been formally proven, this is an intuitively sensible conclusion, since all arrival and departure processes are Poisson of rate $\lambda$, and the distribution of the time spent in a queue is the same regardless of its position in the tandem. If this conjecture is accepted as true, the aforementioned bounds can be refined by adopting the order that yields tighter values. Furthermore, a different order can be selected for each bound as needed.
The following tables present numerical results for tandems consisting of 3 (Table 2), 6 (Table 3) and 10 (Table 4) queues, with a normalized arrival rate $\lambda=1$, equispaced loading rates ranging from $0.1$ to $0.9$ and different orderings. $\rho_{i}$ indicates the loading rate of the $i$-th server in the tandem. Using the simmer library [16] in R [17], 100 simulations are conducted for each server configuration, and the empirical AAoI is computed for each run. The columns age av and age sd show the mean and standard deviation of the AAoI across the 100 runs. In all of the tables, these columns support the conjecture that the AAoI at the end of the tandem appears to be independent of the servers’ ordering. Additionally, in all of the simulations, the empirical AAoI is within the bounds of Equation (4.3.2), collected in the columns age lb and age ub. Likewise, it can also be noted that the interval is narrower when the slowest service rate corresponds to the last server (shaded rows in the tables).
@@
In the especial case of homogeneous servers, $\gamma_{i}=\mu,\ \forall i=1,\ldots,C+1$, (4.3.2) reduces to
@@
that gives and interval of width $2\sqrt{C}/(\mu-\lambda)$. It can be observed that, for faster servers (i.e., larger $\mu$), the interval moves to the left and becomes tighter, while for larger number of servers, the interval moves to the right and gets wider.
4.3.3 M/M/1/1 $\rightarrow$ $\cdot$/M/1$\infty$ tandem
@@
where $Y_{n}$ is the inter-departure time of the second queue, $A_{n}$ is the age of the $n$th update when entering the second queue, and $\Delta^{0}_{HE/M/1/\infty}$ is the AAoI with zero-age updates defined in the following Proposition, whose demonstration is presented in Appendix B.
@@
where $\rho=\mathbb{E}\left[S\right]/\mathbb{E}\left[X\right]=\dfrac{\lambda\gamma}{\mu(\lambda+\gamma)}<1$ and $\sigma$ is the unique root in $(0,1)$ of the equation $\sigma=\tilde{X}(\mu-\mu\sigma)$, with $\tilde{X}(\cdot)$ the Laplace-Stieltjes Transform of the inter-arrival time distribution.
In this tandem, all packets departing from the first queue are received and served in the second queue. Therefore, $A_{n}$ is the system time of the first queue, which is an exponential service time of rate $\gamma$ independent of the inter-departure time of future packets in the second queue. Thus, the second term in Equation (14) reduces to $\mathbb{E}\left[A_{n-1}\right]=1/\gamma$, and the age at the end of this tandem is
@@
4.4 Retrial Queues
A Retrial Queue is a type of queueing model where an arrival from the source —referred to as a primary arrival— temporarily leaves the system when it finds the server busy and returns later. While waiting to re-enter, the packet is said to be in an orbit. When the server becomes available, both the source and the orbit compete for access, each submitting packets based on their own renewal processes. This process is depicted in Figure 6. There are two key differences between a standard queue and the orbit in retrial queues. First, packets in the orbit cannot monitor the server’s state; they can only check their status by attempting to re-enter the system, an event known as a retrial. Furthermore, after a service is completed, there is a period during which the server is idle until a new update arrives, whether it be a primary arrival or a packet from the orbit. Secondly, to better reflect the real-life systems it is intended to represent (customers repeatedly calling a store until they receive service), packets in the orbit generally do not follow any specific order; thus, the service discipline is random.
Figure 6: Basic model for a retrial queue (from [18, Section 3.5]).
Retrial queueing models are generally difficult to analyse, and the model described here is one of the few that can be solved analytically. Therefore, it is not surprising that, to the best of our knowledge, the AoI has not yet been studied in this class of models. However, using the aged-updates scheme, the calculation of AAoI in these models can be significantly simplified, enabling the derivation of either an exact expression or reasonable bounds. For clarity, the specific model considered in this example will first be described, followed by an explanation of how to apply the scheme of aged updates to determine the AAoI for the model.

### v2

@@ 4.1.2 Light traffic
The LT regime occurs when the effective arrival rate approaches zero. Together with the flow conservation principle, this feature is captured by the condition
@@
For a given standard deviation, this implies that the coefficient of variation $\kappa_{Y}$ goes to zero. Consequently, the second term in the right-hand side of Equation 7 and the interval width in Corollary 1 become negligible, which leads to the independent case with very tight bounds.
@@ 4.2 Forwarding
To illustrate this, a zero-wait model is considered, in which the source is aware of the system’s state and can generate a fresh packet as soon as a delivery occurs. This model is idealised, but also insightful in capturing the best-case AoI performance [yates:2021, Section III.B]. In particular, because of the source responsiveness, the time variables in the zero-wait model are identically distributed,
$X_{n+1}=T_{n}=Y_{n}=S_{n}.$ (9)
@@
Figure 3: Error-prone, zero-wait model as a system with aged updates. a) Error-prone, zero-wait standard model. Blue triangles are successful deliveries and generation of a fresh packet, and red triangles ($t_{3}$ and $t_{5}$) mark failed deliveries and subsequent retransmission of the same packet. b) Equivalent error-free, zero-wait model with aged updates. Failed transmissions in the standard model can be thought of as inducing an independent process of initial ages $A_{n}$.
The value of the AAoI in a zero-wait model with error probability $\alpha$, as that depicted in Figure 3a), is given in the following Proposition.
Proposition 1.
The AAoI in a zero-wait, error-prone model with error probability $\alpha$ and exponential service times of rate $\mu$ is
$\Delta_{zw}=\dfrac{2}{\mu(1-\alpha)}.$ (10)
Note that the number of retransmissions for a given packet in this model is a geometric random variable with parameter $\alpha$. Therefore, the effective service time is a geometric sum of exponential times, which results in an exponential distribution of rate $\mu(1-\alpha)$. The result in Proposition 1 then follows directly from [yates:2021, Equation (22)].
Notably, the same expression is obtained by applying Theorem 1 to the error-free model of Figure 3b) with an independent process of initial ages defined by
@@
where $\tilde{S}_{1}\sim\textsl{exp}\left(\mu(1-\alpha)\right)$ and $\tilde{S}_{2}\sim\Gamma\left(2,\mu(1-\alpha)\right)$ are implicitly derived from the number of retransmissions. In Appendix C we provide full details of the proof of Proposition 1 using the aged-updates framework.
@@
$d=2\sigma_{A}=\dfrac{2\sqrt{2\alpha(2-\alpha)}}{\mu(1-\alpha)}.$
However, a more detailed understanding is achieved by examining the bounds directly. To this purpose, Figure 4 plots the correction term and the upper and lower bounds against the error probability $\alpha$, for a fixed service rate $\mu=1$.
Figure 4: Correction term and bounds provided by Corollary 1 in the error-prone, zero-wait model.
The plot reveals some key insights. The bounds are generally tight for low error probabilities, where successful transmissions dominate the system behaviour. As $\alpha$ increases, the upper bound grows considerably, reflecting the high uncertainty and variability introduced by frequent retransmissions. Furthermore, the lower bound becomes negative for $\alpha<2/3$. In this region, the trivial bound $\lambda^{e}\mathbb{E}\left[Y_{n}A_{n-1}\right]\geq 0$ is better for practical purposes. This occurs because the high variability of the initial age $\sigma_{A}$ outweighs its mean value for moderate error rates.
@@ 4.3 Tandem Queues
Some prior works have derived bounds for age metrics in multi-hop systems. For instance, [farazi:2019] establishes fundamental lower bounds on the peak and average age for general multi-source, multi-hop wireless networks under different interference constraints. Similarly, [vikhrova:2022] derives bounds on the average AoI for systems with cross-traffic and various queuing policies. In this section, we demonstrate that our methodology, which relies on a correlation-based decomposition at a single queue, provides a complementary and effective tool for analysing tandem networks.
@@
4.3.1 $\infty \rightarrow M/M/1/\infty$ tandem
This first tandem model serves as a primary illustration of the framework’s ability to handle statistical dependence between the initial ages, which here represents the time spent in the first queue of the tandem, and the inter-departure times in the second queue. This dependence is captured by the explicit computation of the cross-term $\mathbb{E}\left[Y_{n}^{(2}A_{n-1}\right]$ in Lemma 1. First, the theorem that gives the AAoI at the end of this tandem is stated. While this result is already known, this example is useful for confirming that the approach with aged updates can also be accommodated in this model.
Theorem 2 ([kam:2022], Theorem 3111[kam:2022, Equation (13)] contains a typo.).
The AAoI at the end of two infinite-capacity queues in tandem ($\infty \rightarrow M/M/1/\infty$) with $\lambda$-Poisson arrivals, and i.i.d. exponential service times of rates $\gamma$ and $\mu$, is given by
@@ 4.3.1 $\infty \rightarrow M/M/1/\infty$ tandem
[…]
Figure 5: Sample path for age in a tandem of two queues $\infty \rightarrow M/M/1/\infty$. Packets arriving at the second queue have a positive age equal to the system time in the first queue. Note that $Y_{n}^{(1)}$ are inter-departure times in the first queue and inter-arrival times in the second queue.
@@
[…]
@@
Then, Theorem 2 follows by applying Equation 12, Lemma 1 and $\mathbb{E}\left[Y_{n}^{(2)}\right]=1/\lambda$ to Equation 11.
@@
The very dynamics of the system prevent far updates from occurring in this type of tandems. Note that, since in this model $A_{n}=T_{n}^{(1)}$ and $X_{n}=X_{n}^{(2)}=Y_{n}^{(1)}$, the condition for an update to be obsolete is $T_{n}^{(1)}>Y_{n}^{(1)}+T_{n-1}^{(1)}$. However, by the identity (8) and the analogous one for the system time, $T_{n}=S_{n}+\left(T_{n-1}-X_{n}\right)^{+}$, it follows that the condition is equivalent to
$(T_{n-1}^{(1)}-X_{n}^{(1)})^{+}>(X_{n}^{(1)}-T_{n-1}^{(1)})^{+}-T_{n-1}^{(1)},$
which is impossible if either $X_{n}^{(1)}>T_{n-1}^{(1)}$ or $X_{n}^{(1)}<T_{n-1}^{(1)}$.
Remark.
The procedure developed in this section can also derive the AAoI expressions for other well-known tandems. Specifically, the independence between final inter-departure times and initial ages can be exploited in tandems consisting of two single-capacity queues, with both preemption in service [yates:2018, Theorem 2] and without preemption in service [kam:2022, Theorem 1].
[…]
Suppose that the service rate of the $i$-th server is $\gamma_{i},\ i=1,\ldots,C+1$. For consistency with the previous section, the notation $\gamma_{C+1}=\mu$ is used interchangeably to denote the service rate of the last server. Conveniently, the notation $\Delta$ is used to denote the AAoI at the end of the tandem, and $\Delta_{i}$ to denote the AAoI of a single queue with service rate $\gamma_{i}$. By Burke’s Theorem, the departure process at the end of such a tandem remains Poisson of rate $\lambda$, implying that the coefficient of variation of the inter-departure times is $\kappa_{Y}=1$. Additionally, the initial age of the updates entering the last queue is the sum of the system times in each of the preceding queues
@@
$\mathbb{E}\left[A_{n}\right]=\sum\limits_{i=1}^{C}\frac{1}{\gamma_{i}-\lambda}\quad\text{ and }\quad\sigma_{A}=\sqrt{\sum\limits_{i=1}^{C}\frac{1}{(\gamma_{i}-\lambda)^{2}}},\\$
@@
Moreover, evidence suggests the AAoI at the end of this tandem is independent of the ordering of the servers [kam:2022]. Although this hypothesis has not been formally proven, it is intuitively sensible as all arrival and departure processes are Poisson of rate $\lambda$, and the distribution of the time spent in a queue is the same regardless of its position in the tandem. Our numerical simulations strongly support this conjecture, as shown in the following tables. Tables 2, 3 and 4 present results for tandems of 3, 6, and 10 queues respectively, with a normalised arrival rate $\lambda=1$, equispaced loading rates ranging from $0.1$ to $0.9$ and different server orderings. For each configuration, we conducted 100 simulations using the simmer library [ucar:2019] in R [r_core:2022]. The columns age av and age sd show the mean and standard deviation of the AAoI across runs. The consistency of the empirical age across different orderings strongly supports the server-order independence conjecture.
@@
Furthermore, these tables reveal a clear pattern for the bounds from Section 4.3.2: the tightest bounds are consistently achieved when the slowest server is placed last (shaded rows). This occurs because the standard deviation of the initial age $\sigma_{A}$, and therefore the interval width, is minimised when the slowest server is at the end of the tandem. This leads to a significant practical implication: if the server-order independence holds, then the tightest possible bounds for any tandem configuration are given by the bounds calculated for the specific order where the slowest server is last. Therefore, regardless of the actual system configuration, a practitioner can always obtain the most accurate interval estimate by simply assuming the slowest server is in the final position when applying Corollary 1.
In all simulations, Tables 2, 3 and 4 show that the empirical AAoI falls within the theoretical bounds –collected in the columns age lb and age ub–, confirming their validity. However, the bounds can be loose (50-100% error) for unfavourable server orderings, which underscores the importance of the aforementioned ordering strategy to obtain the best possible estimates from our framework.
Furthermore, a closer look at the table reveals that the empirical age lies near the centre of the interval, always closer to the lower bound. This is not coincidental, but a direct consequence of the dependence structure in this tandem; specifically, a systematic negative correlation ($r_{Y,A}<0$) between the initial age $A_{n-1}$ (which is the total system time in previous queues) and the final inter-departure time $Y_{n}$.
The correlation is negative due to the dynamics captured by the Lindley-type equation for the inter-departure times. Expanding the equation of the inter-departure time in the last queue, $Y_{n}^{(C+1)}$, reveals
$Y_{n}^{(C+1)}$ $=T_{n}^{(C+1)}+X_{n}^{(C+1)}-T_{n-1}^{(C+1)}$
$=T_{n}^{(C+1)}+Y_{n}^{(C)}-T_{n-1}^{(C+1)}$
$=T_{n}^{(C+1)}+S_{n}^{(C)}+(X_{n}^{(C)}-T_{n-1}^{(C)})^{+}-T_{n-1}^{(C+1)},$
where $A_{n-1}$ is a function of $T_{n-1}^{(C)}$ and previous system times. When $T_{n-1}^{(C)}$, a component of $A_{n-1}$, is large, the term $(X_{n}^{(C)}-T_{n-1}^{(C)})^{+}$ tends to zero. This suppresses the idle time, making $Y_{n}^{(C+1)}$ smaller. Conversely, when $T_{n-1}^{(C)}$ is small, the idle term is more likely to be positive, increasing $Y_{n}^{(C+1)}$. This inverse relationship, propagated through the tandem, results in the observed negative correlation $r_{Y,A}<0$. For example, using Lemma 1, the covariance between $Y_{n}^{(C+1)}$ and $T_{n-1}^{(C)}$ can be explicitly computed as
$\mathbb{C}\mathrm{ov}{}\left[Y_{n}^{(C+1)},T_{n-1}^{(C)}\right]=\dfrac{1}{\gamma^{2}_{C}}\dfrac{(\gamma_{C}+\gamma_{C+1})(\lambda-\gamma_{C+1})}{\gamma_{C+1}(\gamma_{C}+\gamma_{C+1}-\lambda)},$
which is indeed negative for stable queues ($\gamma_{i}>\lambda$). Since the correction term is $\mathbb{E}\left[A_{n}\right]+r_{Y,A}\cdot\kappa_{Y}\cdot\sigma_{A}$, a negative correlation subtracts from the mean age, pulling the final AAoI down from the upper bound towards the lower bound.
While a precise calculation of $r_{Y,A}$ is complex, its consistent negativity suggests a practical rule of thumb for estimating the correction term in this specific tandem. As the AAoI often lies in the lower half of the bounding interval, a simple yet effective approximation is to use a weighted average leaning towards the lower bound (e.g., $\mathbb{E}\left[Y_{n}A_{n-1}\right]\approx 0.55\cdot LB+0.45\cdot UB$) for a first-order estimate. Likewise, in practice, the upper limit of the interval can be set at $\mathbb{E}\left[A_{n}\right]$.
In the specific case of homogeneous servers, $\gamma_{i}=\mu,\ \forall i=1,\ldots,C+1$, Section 4.3.2 reduces to
@@
that gives and interval of width $2\sqrt{C}/(\mu-\lambda)$. It can be observed that, for faster servers (i.e., larger $\mu$), both limits move to the left and the width becomes tighter, while for a larger number of servers, the interval moves to the right and gets wider.
[…]
We now turn to a system that, to the best of our knowledge, has not yet been analytically characterised from an age perspective: a tandem of two heterogeneous queues. This example will highlight the framework’s ability to derive new analytical results.
@@
where $Y_{n}$ is the inter-departure time of the second queue, $A_{n}$ is the age of the $n$th update when entering the second queue, and $\Delta^{0}_{HE/M/1/\infty}$ is the AAoI with zero-age updates defined in the following Proposition, whose demonstration is presented in Appendix D.
@@
[…]
In this tandem, all packets departing from the first queue are received and served in the second queue. Therefore, $A_{n}$ is the system time of the first queue, which is an exponentially distributed service time of rate $\gamma$ [costa:2016, Section IV.A], and independent of the inter-departure time of future packets in the second queue. Thus, the second term in Equation 14 reduces to $\mathbb{E}\left[A_{n-1}\right]=1/\gamma$, and the age at the end of this tandem results in
@@
Remark.
This tandem also inherently prevents the arrival of far updates. Note that, in this case, $A_{n}=S_{n}^{(1)}$ and $X_{n}=X_{n}^{(2)}=Y_{n}^{(1)}$. Using again identity (8), the condition for an obsolete packet becomes $0>(X_{n}^{(1)}-S_{n-1}^{(1)})^{+}+S_{n-1}^{(1)}$, which is also impossible.
4.4 Retrial queues: A novel application
Retrial queues represent a classic yet complex model in queueing theory. However, their analysis in the context of AoI remains an open challenge. In this section, we leverage the aged-updates framework to provide the first characterisation of the AAoI in an M/M/1 retrial queue.
Figure 6: Basic model for a retrial queue (from [gross:fqt, Section 3.5]).
A retrial queue is a type of queueing model where an arrival from the source —referred to as a primary arrival— temporarily leaves the system when it finds the server busy and returns later. While waiting to re-enter, the packet is said to be in an orbit. When the server becomes available, both the source and the orbit compete for access, each submitting packets based on its own arrival process. The system is sketched in Figure 6. There are two key differences between a standard queue and the orbit in retrial queues. First, packets in the orbit cannot monitor the server’s state; they can only check their status by attempting to re-enter the system, an event known as a retrial. Furthermore, after a service is completed, there is a period during which the server is idle until a new update arrives, whether it be a primary arrival or a packet from the orbit. Secondly, to better reflect the real-life systems it is intended to represent (such as customers repeatedly calling a store until they receive service), packets in the orbit generally do not follow any specific order; thus, the service discipline is random.
Retrial queueing models are generally difficult to analyse, and the model described here is one of the few that can be solved analytically. Therefore, it is not surprising that, to the best of our knowledge, the AoI has not yet been studied in this class of models. However, using the aged-updates scheme, the calculation of AAoI can be significantly simplified, enabling the derivation of either an exact expression or reasonable bounds. For clarity, the specific model considered in this example will first be described, followed by an explanation of how to apply the scheme of aged updates to determine the AAoI.